Education Level Male Female Total
High School or Less 2136 2409 4545
Bachelor's Degree 2827 2854 5681
Master's Degree 385 352 737
Ph.D. 69 57 126
Total 5417 5672 11,089

A person is selected at random. Compute the following probabilities.
(a) What is the probability that the selected person does not have a Ph.D.?

(b) What is the probability that the selected person does not have a Master's degree?

(c) What is the probability that the selected person is female or has a Master's degree?

(d) What is the probability that the selected person is male or has a Ph.D.?

Answer by ewatrrr(24785)


Hi
Education Level Male Female Total
High School or Less 2136 2409 4545
Bachelor's Degree 2827 2854 5681
Master's Degree 385 352 737
Ph.D. 69 57 126
Total 5417 5672 11,089
a) P(NOT PhD) = 1 - P(PhD) = 1 - 126/11,089
b) P(NOT Masters) = 1 - P(M) = 1 - 737/11,089
c) P(F 0r Masters) = (F) + (Masters) - (Both) = 5672/11089 + 737/11089 - 352/11089
d) P(M 0r PhD) = (1/11089) [(5417) + (126) -(57)]